The accounting department of your company has just delivered a draft of the current years financial statements to you. The summary is as folows of the Year End of the Year 550,000 210,000 40,000 554,000 205,000 49,000 106,800 Liabilises income for the Year Shares Outstanding 22.000 22,000 You discovered that they have not adjusted for estimated bad debt expenses of $9,800. For each of the following ratios, calculate 1. The ratio that would have resulted had the error not been discovered (i.e the incorrect ratio) 2. The correct ratio Incorrect Correct 2 ROA ROE Debt Ratio S EPS 10
Answer #1

First we will do without including badbedt

ROA = net income / average assets

= 106800/(550000+554000)/2= 0.1927= 19.27%

ROE= net income /avg shareholders equity

= 106800/(340000+349000)/2=31%

Debt ratio over the year= avgdebt/average assets

= (210000+205000)/550000+554000= 37.5%

EPS = net income/ total outstanding shares

106800/22000= 4.85

In 2nd case we should create provision for badbedt which decreases assets by 9800 and net income by 9800 there by liabilities decreased by 9800

New total asset 544200

New total liabilities 195200

New netincome 97000

1) ROA = 97000/(550000+544200)/2= 17.73%

2)ROE= 97000/(340000+ 339200)/2= 28.56%

3)DEBT RATIO=(210000+195200)/(550000+544200)= 37%

4)EPS= 97000/22000= 4.4